A shock absorber acts like an elastic spring of constant k. It compresses by 2.5 cm when a mass m is placed on it. When the absorber-mass system is pressed and left, it begins to oscillate. Calculate the frequency of oscillations. Take g : 10 ms^-2
1
Expert's answer
2019-05-08T09:42:26-0400
"f=\frac{1}{2\pi}\sqrt{\frac{k}{m}}"

"kx=mg"

"f=\frac{1}{2\pi}\sqrt{\frac{g}{x}} =\frac{1}{2\pi}\sqrt{\frac{10}{0.025}} =3.2 \ Hz"